OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. problem to come...

OpenStudy (anonymous):

\[f(x)=(x-7)/(x+3) and g(x)=(-3x-7)/(x-1)\]

OpenStudy (anonymous):

the first step i know is to write.... im working on it...

OpenStudy (zzr0ck3r):

yeah its ugly

OpenStudy (anonymous):

\[((-3x-7)/(x-1)-7)/ ((-3-7)/(x-1)+3)\]

OpenStudy (anonymous):

i get that...but i dont know hw to simplify it all

OpenStudy (zzr0ck3r):

is that f(g(x)) or g(f(x))?

OpenStudy (anonymous):

thats f(g(x))

OpenStudy (zzr0ck3r):

\[f(x)=\frac{x-7}{x+3}\\g(x) = \frac{-3x-7}{x-1}\\ f(g(x)) =\frac{g(x)-7}{g(x)+3}=\frac{\frac{-3x-7}{x-1}-7}{\frac{-3x-7}{x-1}+3} \] is this what you did?

OpenStudy (anonymous):

Yep

OpenStudy (zzr0ck3r):

ok \[\frac{\frac{-3x-7}{x-1}-7}{\frac{-3x-7}{x-1}+3}\] multiply everything by (x-1) \[\frac{(x-1)\frac{-3x-7}{x-1}-7(x-1)}{(x-1)\frac{-3x-7}{x-1}+(x-1)3}=\\\frac{-3x-7-7x+7}{-3x-7+3x-3}=\frac{-10x}{-10}=x\]

OpenStudy (anonymous):

oh boy...tiny writing

OpenStudy (zzr0ck3r):

mac or pc?

OpenStudy (anonymous):

pc XD but i got it

OpenStudy (zzr0ck3r):

do the steps make sense?

OpenStudy (anonymous):

completely :D now can you help me with g(f(x))?

OpenStudy (zzr0ck3r):

\[g(f(x))= \frac{-3f(x)-7}{f(x)-1}=\frac{-3\frac{x-7}{x+3}-7}{\frac{x-7}{x+3}-1}\]

OpenStudy (zzr0ck3r):

multiply every term by (x+3) and tell me what you have

OpenStudy (anonymous):

ok :) let me try :)

OpenStudy (anonymous):

not the right thing.. thats for sure....

OpenStudy (zzr0ck3r):

\[\frac{-3(x-7)-7x-21}{x-7-x-3}=\frac{-10x}{-10}=x\]

OpenStudy (zzr0ck3r):

this is after you multiply everything by (x+3)

OpenStudy (zzr0ck3r):

I hope you understand this, because they were exactly alike, and you should be able to do it now. Unless you don't understand something in which case tell me

OpenStudy (anonymous):

how did you get the numerator to that...becuse it cancels out its denominator but where did the -7x and 21 come from

OpenStudy (anonymous):

NEVERMIND

OpenStudy (zzr0ck3r):

\[\frac{-3\frac{x-7}{x+3}(x+3)-7(x+3)}{\frac{x-7}{x+3}(x+3)-1(x+3)}=\frac{-3(x-7)-7x-21}{x-7-x-3}=\frac{-3x+21-7x-21}{x-7-x-3}\]

OpenStudy (zzr0ck3r):

ok:)

OpenStudy (anonymous):

Thank you!

OpenStudy (zzr0ck3r):

np